OpenStudy (anonymous):

NEED HELP WITH 2 MATH QUESTIONS PLZ!

OpenStudy (anonymous):

What

OpenStudy (anonymous):

tell me them

OpenStudy (displayerror):

For question 1:The total probability should always add up to 1 and the probability of an event can never be negative. For question 2: Just add the probabilities. You're given a graph of the probability distribution and the range of values that you have to calculate the probability for. For question 3: The expected value is given by the sum of X times P: \[\sum\left( X \times P \right)\]

OpenStudy (anonymous):

can you work it out for me for number 1 and 3

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

NEED HELP WITH 1 AND 3 PLZ

OpenStudy (anonymous):

@AnswerMyQuestions